# Quantitative Propagation of Chaos for 2D Viscous Vortex Model with General Circulations on the Whole Space

### v2

@@ 2.1. Time Evolution of Relative Entropy
$\displaystyle$=\,= $\displaystyle\frac{1}{N}\int_{\mathbb{D}^{N}}\partial_{t}G^{N}(\log G^{N}+1-\log g^{N})+\frac{G^{N}}{g^{N}}\partial_{t}g^{N}\,\mathrm{d}Z^{N}$
$\displaystyle$=\,= $\displaystyle\frac{1}{N}\int_{\mathbb{D}^{N}}\Big{(}\sigma\sum_{i=1}^{N}\Delta_{x_{i}}G^{N}-\frac{1}{N}\sum_{i,j=1}^{N}m_{j}K(x_{i}-x_{j})\cdot\nabla_{x_{i}}G^{N}\Big{)}(\log G^{N}+1-\log g^{N})\,\mathrm{d}Z^{N}$
@@
$\displaystyle$== $\displaystyle-\frac{\sigma}{k}\sum_{i=1}^{k}\int_{\mathbb{D}^{k}}G^{N,k}\Big{|}\nabla_{x_{i}}\log\frac{G^{N,k}}{g^{k}}\Big{|}^{2}\,\mathrm{d}Z^{k}$
@@
$\displaystyle$=\,= $\displaystyle\frac{1}{k}\int_{\mathbb{D}^{k}}\partial_{t}G^{N,k}(\log G^{N,k}+1-\log g^{k})+\frac{G^{N,k}}{g^{k}}\partial_{t}g^{k}\,\mathrm{d}Z^{k}$
$\displaystyle$=\,= $\displaystyle\frac{1}{k}\int_{\mathbb{D}^{k}}\Big{(}\sigma\sum_{i=1}^{k}\Delta_{x_{i}}G^{N,k}-\frac{1}{N}\sum_{i,j=1}^{k}m_{j}K(x_{i}-x_{j})\cdot\nabla_{x_{i}}G^{N,k}$
@@ 2.2. Law of Large Numbers and Large Deviation Type Theorems
×$\int_{{\mathbb{R}}\times\mathbb{R}^{2}}\phi(z,w)\bar{\rho}(z)\,\mathrm{d}z=0,\,\mbox{for any\, }w,\quad\int_{{\mathbb{R}}\times\mathbb{R}^{2}}\phi(z,w)\bar{\rho}(w)\,\mathrm{d}w=0,\,\mbox{for any\, }z,$ (2.2)
@@
(×$\int_{({\mathbb{R}}\times\mathbb{R}^{2})^{N}}\bar{\rho}^{\otimes N}\exp\Big{(}\frac{1}{N}\sum_{i,j=1}^{N}\phi(z_{i},z_{j})\Big{)}\,\mathrm{d}Z^{N}\leq C<\infty.$
@@
×$\int_{{\mathbb{R}}\times\mathbb{R}^{2}}\phi(z,w)\bar{\rho}(z)\,\mathrm{d}z=0,\,\mbox{for any\, }w,\quad\int_{{\mathbb{R}}\times\mathbb{R}^{2}}\phi(z,w)\bar{\rho}(w)\,\mathrm{d}w=0,\,\mbox{for any\, }z,$ (2.5)
@@
Finally we present two lemmas, say the Donsker-Varadhan inequality for the variational expression for the relative entropy, and the chain rule of the relative entropy for disintegrated probability measures.
@@
$\int_{\mathbb{D}^{N}}\Phi\rho_{N}\,\mathrm{d}Z^{N}\leq\frac{1}{\eta}H_{N}(\rho_{N}|\bar{\rho}^{N})+\frac{1}{\eta N}\log\int_{\mathbb{D}^{N}}\exp(N\eta\Phi)\,\mathrm{d}Z^{N}.$
Lemma 2.4 (Chain Rule, borrowed from [34]).
@@
We refer the proof of Lemma 2.3 to [30] and the proof of Lemma 2.4 to Budhiraja-Dupuis [8].
@@ 2.3. Global Relative Entropy Control
$-\frac{1}{N^{2}}\sum_{i,j=1}^{N}\int_{\mathbb{D}^{N}}G^{N}\Big{(}m_{j}K(x_{i}-x_{j})-K\ast\omega(x_{i})\Big{)}\cdot\nabla\log g(m_{i},x_{i})\,\mathrm{d}Z^{N}.$
@@
$\displaystyle$++ $\displaystyle\,\frac{1}{N^{2}}\sum_{i,j=1}^{N}\int_{\mathbb{D}^{N}}G^{N}\Big{(}m_{j}V(x_{i}-x_{j})-V\ast\omega(x_{i})\Big{)}:\frac{\nabla_{x_{i}}^{2}g^{N}}{g^{N}}\,\mathrm{d}Z^{N}.$ (2.8)
@@
[…]
@@
$\displaystyle$++ $\displaystyle\frac{2}{N\sigma}\sum_{i=1}^{N}\int_{\mathbb{D}^{N}}G^{N}\Big{|}\frac{1}{N}\sum_{j=1}^{N}\Big{(}m_{j}V(x_{i}-x_{j})-V\ast\omega(x_{i})\Big{)}\Big{|}^{2}|\nabla\log g(m_{i},x_{i})|^{2}\,\mathrm{d}Z^{N}.$ (2.10)
@@
$\displaystyle$++ $\displaystyle\frac{2}{N^{2}}\int_{\mathbb{D}^{N}}G^{N}|V\ast\omega(x_{1})|\Big{|}\sum_{j=2}^{N}\Big{(}m_{j}V(x_{1}-x_{j})-V\ast\omega(x_{1})\Big{)}\Big{|}|\nabla\log g(m_{1},x_{1})|^{2}\,\mathrm{d}Z^{N}$
$\displaystyle$++ $\displaystyle\int_{\mathbb{D}^{N}}G^{N}\Big{|}\frac{1}{N}\sum_{j=2}^{N}\Big{(}m_{j}V(x_{1}-x_{j})-V\ast\omega(x_{1})\Big{)}\Big{|}^{2}|\nabla\log g(m_{1},x_{1})|^{2}\,\mathrm{d}Z^{N}.$
@@
[…]
@@
[…]
$\displaystyle$++ $\displaystyle\,\int_{\mathbb{D}^{N}}\Big{|}\frac{1}{N}\sum_{j=2}^{N}\Big{(}m_{j}V(x_{1})-{\mathbb{E}}{\mathcal{M}}\cdot V(x_{1})\Big{)}\Big{|}^{2}|\nabla\log g(m_{1},x_{1})|^{2}G^{N}\,\mathrm{d}Z^{N}.$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\displaystyle$++ $\displaystyle\,\frac{C}{1+t}\int_{\mathbb{D}^{N-1}}H_{1}(P_{z_{2},\cdots,z_{N}}^{N,1}|g)G^{N,N-1}\,\mathrm{d}z_{2}\cdots\,\mathrm{d}z_{N}.$
@@
The Grönwall’s inequality will give our desired bound for the global relative entropy. Applying the sub-additivity property and the CKP inequality as in [52], we resolve the $L^{1}$ strong propagation of chaos.
@@ 2.4. Local Relative Entropy Control
[…]
@@
$\displaystyle$++ $\displaystyle\frac{1}{kN}\sum_{i,j=1}^{k}\int_{\mathbb{D}^{k}}G^{N,k}\Big{(}m_{j}V(x_{i}-x_{j})-V\ast\omega(x_{i})\Big{)}:\frac{\nabla_{x_{i}}^{2}g^{k}}{g^{k}}\,\mathrm{d}Z^{k}.$
@@
$\displaystyle$== $\displaystyle\,\frac{\sigma}{4}I_{k}(G^{N,k}|g^{k})+\frac{C}{1+t}\Big{(}H_{k}(G^{N,k}|g^{k})+\frac{k}{N^{2}}\Big{)},$
@@
$\displaystyle$++ $\displaystyle\frac{1}{\sigma}\int_{\mathbb{D}^{k+1}}G^{N,k}\Big{|}m_{k+1}K(x_{1}-x_{k+1})\frac{G^{N,k+1}}{G^{N,k}}-K\ast\omega(x_{1})\Big{|}^{2}\,\mathrm{d}Z^{k+1}.$
@@
[…]
@@
[…]
@@
Inspired by [55], we introduce the following new ODE hierarchy proposition to deal with the evolution of the local relative entropy, as an extension to more general growth-in-time functions beyond the previous work [55].
@@
Let $T>0$ and let $x_{k},y_{k}\in C^{1}([0,T];[0,\infty))$ for any $1\leq k\leq N$. Suppose that $x_{k+1}\geq x_{k}$ for all $1\displaystyle 1\leq k\leq N-$ 1. Assume that there exists real numbers $c_{1}>c_{2}\geq 0$ and some non-negative time decay function $h(t)$ such that for all $t\in[0,T]$,
@@
This will conclude our proof of the main results.

### v3

@@ 2.1. Time Evolution of Relative Entropy
$\displaystyle=\,$ $\displaystyle\frac{1}{N}\int_{\mathbb{D}^{N}}\partial_{t}G^{N}(\log G^{N}+1-\log g^{N})+\frac{G^{N}}{g^{N}}\partial_{t}g^{N}\,\mathrm{d}Z^{N}$
$\displaystyle=\,$ $\displaystyle\frac{1}{N}\int_{\mathbb{D}^{N}}\Big(\sigma\sum_{i=1}^{N}\Delta_{x_{i}}G^{N}-\frac{1}{N}\sum_{i,j=1}^{N}m_{j}K(x_{i}-x_{j})\cdot\nabla_{x_{i}}G^{N}\Big)(\log G^{N}+1-\log g^{N})\,\mathrm{d}Z^{N}$
@@
$\displaystyle=$ $\displaystyle-\frac{\sigma}{k}\sum_{i=1}^{k}\int_{\mathbb{D}^{k}}G^{N,k}\Big|\nabla_{x_{i}}\log\frac{G^{N,k}}{g^{k}}\Big|^{2}\,\mathrm{d}Z^{k}$
@@
$\displaystyle=\,$ $\displaystyle\frac{1}{k}\int_{\mathbb{D}^{k}}\partial_{t}G^{N,k}(\log G^{N,k}+1-\log g^{k})+\frac{G^{N,k}}{g^{k}}\partial_{t}g^{k}\,\mathrm{d}Z^{k}$
$\displaystyle=\,$ $\displaystyle\frac{1}{k}\int_{\mathbb{D}^{k}}\Big(\sigma\sum_{i=1}^{k}\Delta_{x_{i}}G^{N,k}-\frac{1}{N}\sum_{i,j=1}^{k}m_{j}K(x_{i}-x_{j})\cdot\nabla_{x_{i}}G^{N,k}$
@@ 2.2. Law of Large Numbers and Large Deviation Type Theorems
$\int_{{\mathbb{R}}\times\mathbb{R}^{2}}\phi(z,w)\bar{\rho}(z)\,\mathrm{d}z=0,\,\mbox{for any\, }w,\quad\int_{{\mathbb{R}}\times\mathbb{R}^{2}}\phi(z,w)\bar{\rho}(w)\,\mathrm{d}w=0,\,\mbox{for any\, }z,$ (2.2)
@@
$\int_{({\mathbb{R}}\times\mathbb{R}^{2})^{N}}\bar{\rho}^{\otimes N}\exp\Big(\frac{1}{N}\sum_{i,j=1}^{N}\phi(z_{i},z_{j})\Big)\,\mathrm{d}Z^{N}\leq C<\infty.$
@@
$\int_{{\mathbb{R}}\times\mathbb{R}^{2}}\phi(z,w)\bar{\rho}(z)\,\mathrm{d}z=0,\,\mbox{for any\, }w,\quad\int_{{\mathbb{R}}\times\mathbb{R}^{2}}\phi(z,w)\bar{\rho}(w)\,\mathrm{d}w=0,\,\mbox{for any\, }z,$ (2.5)
@@
Finally, we present two technical lemmas that shall be used in our proof. The first one is the famous Donsker-Varadhan inequality, which comes from a variational formula of the relative entropy. This inequality allows us to control the expectation of a test function with respect to a non-factorized density $\rho_{N}$ by the expectation of another exponential test function with respect to a factorized density $\bar{\rho}^{\otimes N}$ and the relative entropy between the two densities. The expectation with respect to a factorized density can be estimated by the Law of Large Numbers Theorem 2.3 and the Large Deviation Theorem 2.4 above.
@@
$\int_{\mathbb{D}^{N}}\Phi\rho_{N}\,\mathrm{d}Z^{N}\leq\frac{1}{\eta}H_{N}(\rho_{N}|\bar{\rho}^{\otimes N})+\frac{1}{\eta N}\log\int_{\mathbb{D}^{N}}\exp(N\eta\Phi)\bar{\rho}^{\otimes N}\,\mathrm{d}Z^{N}.$
The proof of Lemma 2.3 is rather standard in the literature. We refer to [30, Lemma 1] for instance. The second one is the chain rule of the relative entropy between two disintegrated probability measures, which allows us to estimate the relative entropy between two conditional measures.
Lemma 2.4 (Chain Rule of Relative Entropy).
@@
We refer the proof of Lemma 2.4 to Budhiraja-Dupuis [8].
@@ 2.3. Global Relative Entropy Control
$\frac{1}{N^{2}}\sum_{i,j=1}^{N}\int_{\mathbb{D}^{N}}G^{N}\Big(m_{j}K(x_{i}-x_{j})-K\ast\omega(x_{i})\Big)\cdot\nabla\log g(m_{i},x_{i})\,\mathrm{d}Z^{N}.$
@@
$\displaystyle+$ $\displaystyle\,\frac{1}{N^{2}}\sum_{i,j=1}^{N}\int_{\mathbb{D}^{N}}G^{N}\Big(m_{j}V(x_{i}-x_{j})-V\ast\omega(x_{i})\Big):\frac{\nabla_{x_{i}}^{2}g^{N}}{g^{N}}\,\mathrm{d}Z^{N}.$ (2.8)
@@
$\displaystyle\leq$ $\displaystyle\,\frac{2CB_{1}}{1+t}\int_{\mathbb{D}}\exp\Big(\frac{B_{2}}{2}-\frac{1}{2}\log(1+t)-\frac{1}{2}\log g(z)\Big)g(z)\,\mathrm{d}z$
$\displaystyle\leq$ $\displaystyle\,\frac{C}{1+t}\int_{\mathbb{D}}\frac{1}{\sqrt{1+t}}\sqrt{g(z)}\,\mathrm{d}z\leq\frac{C}{1+t},$
@@
$\displaystyle+$ $\displaystyle\frac{2}{N\sigma}\sum_{i=1}^{N}\int_{\mathbb{D}^{N}}G^{N}\Big|\frac{1}{N}\sum_{j=1}^{N}\Big(m_{j}V(x_{i}-x_{j})-V\ast\omega(x_{i})\Big)\Big|^{2}|\nabla\log g(m_{i},x_{i})|^{2}\,\mathrm{d}Z^{N}.$ (2.10)
@@
$\displaystyle+$ $\displaystyle\frac{2}{N^{2}}\int_{\mathbb{D}^{N}}G^{N}|V\ast\omega(x_{1})|\Big|\sum_{j=2}^{N}\Big(m_{j}V(x_{1}-x_{j})-V\ast\omega(x_{1})\Big)\Big||\nabla\log g(m_{1},x_{1})|^{2}\,\mathrm{d}Z^{N}$
$\displaystyle+$ $\displaystyle\int_{\mathbb{D}^{N}}G^{N}\Big|\frac{1}{N}\sum_{j=2}^{N}\Big(m_{j}V(x_{1}-x_{j})-V\ast\omega(x_{1})\Big)\Big|^{2}|\nabla\log g(m_{1},x_{1})|^{2}\,\mathrm{d}Z^{N}.$
@@
$\displaystyle\leq$ $\displaystyle\,\frac{C}{N}\Big(\frac{1}{\eta}H_{1}(G^{N,1}|g)+\frac{1}{\eta}\log\int_{\mathbb{D}}\exp\Big(\frac{\eta B_{1}}{1+t}(B_{2}-\log(1+t)-\log g(z))\Big)g(z)\,\mathrm{d}z\Big)$
$\displaystyle\leq$ $\displaystyle\,\frac{C}{1+t}\Big(\frac{1}{N}H_{N}(G^{N}|g^{N})+\frac{1}{N}\log\int_{\mathbb{D}}\frac{C}{\sqrt{1+t}}\sqrt{g(z)}\,\mathrm{d}z\Big)$
$\displaystyle\leq$ $\displaystyle\,\frac{C}{1+t}\Big(H_{N}(G^{N}|g^{N})+\frac{1}{N}\Big).$
@@
$\displaystyle\triangleq$ $\displaystyle\,\int_{\mathbb{D}^{N}}\Big|\frac{1}{N}\sum_{j=2}^{N}\Big(m_{j}V(x_{1}-x_{j})-m_{j}V(x_{1})+{\mathbb{E}}{\mathcal{M}}\cdot V(x_{1})-V\ast\omega(x_{1})\Big)\Big|^{2}|\nabla\log g(m_{1},x_{1})|^{2}G^{N}\,\mathrm{d}Z^{N}$
$\displaystyle+$ $\displaystyle\,\int_{\mathbb{D}^{N}}\Big|\frac{1}{N}\sum_{j=2}^{N}\Big(m_{j}V(x_{1})-{\mathbb{E}}{\mathcal{M}}\cdot V(x_{1})\Big)\Big|^{2}|\nabla\log g(m_{1},x_{1})|^{2}G^{N}\,\mathrm{d}Z^{N}.$
@@
$\displaystyle\leq$ $\displaystyle\,C\sup_{m_{1},|x_{1}|\leq 2|x_{j}|}|\nabla\log g(m_{1},x_{1})|^{2}$
$\displaystyle\leq$ $\displaystyle\,\frac{C}{1+t}\Big(1+\frac{|x_{j}|^{2}}{1+t}\Big),$
@@
$\displaystyle\leq$ $\displaystyle\,C+C\int_{|x_{1}|\geq 2|y|}mg(m,y)\frac{|y|}{|x_{1}|}\,\mathrm{d}m\,\mathrm{d}y$
$\displaystyle\leq$ $\displaystyle\,C\Big(1+\frac{\sqrt{1+t}}{|x_{1}|}\Big),$
@@
$\displaystyle\leq$ $\displaystyle\,\frac{C}{1+t}+C\Big(1+\frac{|x_{j}|^{2}}{|x_{1}|^{2}}+\frac{1+t}{|x_{1}|^{2}}\Big)\frac{|x_{1}|^{2}}{(1+t)^{2}}$
$\displaystyle\leq$ $\displaystyle\,\frac{C}{1+t}\Big(1+\frac{|x_{j}|^{2}}{1+t}\Big).$
@@
$\displaystyle\leq$ $\displaystyle\,\frac{1}{1+t}+\frac{1}{\eta}H_{1}(P_{z_{2},\cdots,z_{N}}^{N,1}|g)+\frac{1}{\eta}\log\int_{\mathbb{D}}\exp\Big(\frac{\eta|x_{1}|^{2}}{(1+t)^{2}}\Big)g(m_{1},x_{1})\,\mathrm{d}z_{1}$
$\displaystyle\leq$ $\displaystyle\,\frac{C}{1+t}H_{1}(P_{z_{2},\cdots,z_{N}}^{N,1}|g)+\frac{C}{1+t}.$
@@
$\displaystyle+$ $\displaystyle\,\frac{C}{1+t}\int_{\mathbb{D}^{N-1}}H_{1}(P_{z_{2},\cdots,z_{N}}^{N,1}|g)G^{N,N-1}\,\mathrm{d}z_{2}\cdots\,\mathrm{d}z_{N}.$
@@
By the Grönwall inequality, we obtain an upper bound for the global relative entropy $H_{N}(G^{N}|g^{N})$ from the above differential inequality:
$H_{N}(G^{N}|g^{N})\leq(1+t)^{C}\Big(H_{N}(G_{0}^{N}|g_{0}^{N})+\frac{1}{N}\Big).$
This proves Theorem 2.7. By the sub-additivity property of the relative entropy and the condition on the initial data (see Remark 1.3), we have
$H_{k}(G^{N,k}|g^{k})\leq H_{N}(G^{N}|g^{N})\leq\frac{1}{N}(1+t)^{C}.$
Using the CKP inequality, we have the control of the $L^{1}$ distance:
$\|\omega^{N,k}-\omega^{\otimes k}\|_{L^{1}({\mathbb{R}}^{2k})}\leq\|G^{N,k}-g^{k}\|_{L^{1}(\mathbb{D}^{k})}\leq\sqrt{2kH_{k}(G^{N,k}|g^{k})}\leq\frac{\sqrt{k}}{\sqrt{N}}M(1+t)^{M},$
which completes the proof of Corollary 1.1.
@@ 2.4. Local Relative Entropy Control
$\displaystyle\triangleq$ $\displaystyle-\frac{1}{k}\sum_{i=1}^{k}\frac{1}{N}\sum_{j=1}^{k}\int_{\mathbb{D}^{k}}G^{N,k}\Big(m_{j}K(x_{i}-x_{j})-K\ast\omega(x_{i})\Big)\cdot\nabla\log g(m_{i},x_{i})\,\mathrm{d}Z^{k}$
@@
$\displaystyle+$ $\displaystyle\frac{1}{kN}\sum_{i,j=1}^{k}\int_{\mathbb{D}^{k}}G^{N,k}\Big(m_{j}V(x_{i}-x_{j})-V\ast\omega(x_{i})\Big):\frac{\nabla_{x_{i}}^{2}g^{k}}{g^{k}}\,\mathrm{d}Z^{k}.$
@@
$\displaystyle=$ $\displaystyle\,\frac{\sigma}{4}I_{k}(G^{N,k}|g^{k})+\frac{C}{1+t}\Big(H_{k}(G^{N,k}|g^{k})+\frac{k}{N^{2}}\Big),$
@@
$\displaystyle+$ $\displaystyle\frac{1}{\sigma}\int_{\mathbb{D}^{k+1}}G^{N,k}\Big|m_{k+1}K(x_{1}-x_{k+1})\frac{G^{N,k+1}}{G^{N,k}}-K\ast\omega(x_{1})\Big|^{2}\,\mathrm{d}Z^{k+1}.$
@@
$\displaystyle\leq$ $\displaystyle\,\int|V|\Big|\frac{\nabla m_{1}}{m_{1}}-\frac{\nabla m_{2}}{m_{2}}\Big|\,\mathrm{d}m_{1}+\Big|\int V\frac{\nabla m_{2}}{m_{2}}\,\mathrm{d}(m_{1}-m_{2})\Big|$
$\displaystyle\leq$ $\displaystyle\,\|V\|_{L^{\infty}}\sqrt{\int\Big|\nabla\log\frac{m_{1}}{m_{2}}\Big|^{2}\,\mathrm{d}m_{1}}+\lambda^{-1}|\langle m_{1}-m_{2},\lambda V\nabla\log m_{2}\rangle|$
$\displaystyle\leq$ $\displaystyle\,\|V\|_{L^{\infty}}\sqrt{I(m_{1}|m_{2})}+\lambda^{-1}\sqrt{2\Big(1+\log\int e^{\lambda^{2}|V|^{2}|\nabla\log m_{2}|^{2}}\,\mathrm{d}m_{2}\Big)H(m_{1}|m_{2})}.$
@@
$\displaystyle\leq$ $\displaystyle\,\frac{1}{\sigma}\int_{\mathbb{D}^{k}}G^{N,k}\Big[A^{2}\|V\|_{L^{\infty}}^{2}I\Big(\frac{G^{N,k+1}}{G^{N,k}}\Big|g\Big)$
@@
Inspired by [55], we introduce the following new ODE hierarchy proposition to deal with the evolution of the local relative entropy.
@@
Let $T>0$ and let $x_{k},y_{k}\in C^{1}([0,T];[0,\infty))$ for any $1\leq k\leq N$. Suppose that $x_{k+1}\geq x_{k}$ for all $1\leq k\leq N-1$. Assume that there exists real numbers $c_{1}>c_{2}\geq 0$ and some non-negative time decay function $h(t)$ such that for all $t\in[0,T]$,
@@
Using the CKP inequality, we have the control of the $L^{1}$ distance:
$\|\omega^{N,k}-\omega^{\otimes k}\|_{L^{1}({\mathbb{R}}^{2k})}\leq\|G^{N,k}-g^{k}\|_{L^{1}(\mathbb{D}^{k})}\leq\sqrt{2kH_{k}(G^{N,k}|g^{k})}\leq\frac{k}{N}M(1+t)^{M},$
which completes the proof of Corollary 1.2.